OpenStudy (anonymous):

(8.76 x 109)(6.52 x 10-3)/13.27 x 105

OpenStudy (igreen):

You can't just use a calculator and simplify then divide?

OpenStudy (anonymous):

the underlined word is beautiful

OpenStudy (anonymous):

it should be something like 31.28

OpenStudy (igreen):

\(\dfrac{(8.76 \times 109)(6.52 \times 10-3)}{13.27 \times 105}\) \(\dfrac{(954.84)(6.52 \times 7)}{1393.35}\) \(\dfrac{(954.84)(45.64)}{1393.35}\) \(\dfrac{43578.8976}{1393.35}\) \(31.2763466\)

OpenStudy (igreen):

@caesarjua You shouldn't just give out the answer without any explanation or steps.

OpenStudy (anonymous):

hey man, i'm just trying to help

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Jaylin081414

OpenStudy (igreen):

Yes but it's against the sites rules to give out direct answers without any explanation. Read more here: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i didn't even know there were rules

OpenStudy (brooke2353):

469936.702336